# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ 4 Results: PINN for inverse problem and transfer learning
In the previous sections, we discussed the results of pre-trained neural networks for empirical formulae and DeepONets for EGR and VGT states. We introduced a hybrid method that incorporates these pre-trained networks within the PINN framework in section 2.4 and hybrid multi-stage transfer learning in section 2.5. In this section, we present results for the inverse problem solved using the hybrid multi-stage transfer learning approach outlined in section 2.5. We consider two scenarios: i) training with clean data, and ii) training with noisy data. Since we use simulated data for model training, we added Gaussian noise to these simulated outputs to mimic real-world conditions. Based on expected noise from field sensors, we added Gaussian noise values of 3%, 3%, 1%, 5% and 3% to simulated outputs for $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ respectively. The input signal trains required for the Simulink model viz. control input $u=\{u_{\delta},u_{egr},u_{vgt}\}$ and engine speed $n_{e}$, are obtained from field measurements on an engine test bed. As discussed in section 2.5, we train the hybrid PINN model on individual one minute data segments. In the first segment, the complete hybrid PINN model is trained. Complete training of this hybrid PINN model is computationally expensive since it requires learning all network parameters and requires more training epochs. To reduce computational time, we transfer the network weights and biases to the multi-stage transfer learning model for subsequent time segments. Refer Table 10 showing the improvement in computational performance with the transfer learning approach. For each minute of data segment, 301 residual points were considered in the physics-informed loss term of equation 21. Heuristically, we identified that with clean data, 151 residual points were sufficient for accurate parameter and state estimations. However, for noisy data, 301 residual points were used during the training process. The data sampling frequency was 5 Hz.
As discussed earlier in section 2.4, we approximate $p_{im}$, $p_{em}$, $\omega_{t}$, $T_{1}$ and $x_{r}$ using neural networks. We also incorporate mask function and scaling for the unknown parameters, as shown in Table 1. We note that output transformation, mask function, and scaling are important to ensure the network outputs and unknown parameters lie in $\mathbb{R^{+}}$. Additionally, we incorporate self-adaptive weights for the data loss term of $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$, initialized from a Gaussian random variable with mean $\dfrac{\alpha X_{data}}{\max(X_{data})}$, $X=\{p_{em},p_{im},\omega_{t},W_{egr}\}$ and standard deviation $\sigma=0.1$, where $\alpha$ is a heuristically determined weight multiplier ($10^{3}$, $10^{3}$, $10^{3}$, and $2\times 10^{3}$ for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ respectively).
@@
The self-adaptive weight $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ are initialized from a Gaussian random variable with mean $1500$ and standard deviation 0.1. A softplus masking function is applied to the self-adaptive weights to ensure that they remain positive. The unknown parameters are initialized from a Gaussian random variable with standard deviation $\sigma=0.2$ and mean near the true values of the parameters. The parameters of the neural network, unknown engine parameters, and self-adaptive weights are optimized using the Adam optimizer [31] with piece-wise constant learning rates. To estimate the epistemic uncertainty, we train the hybrid PINN model separately for each data segment to generate an ensemble of 30 runs. This uncertainty arises from the random initialization of self-adaptive weights $\mathbf{\lambda}$, unknown parameter values $\mathbf{\Lambda}$, and dropouts included in the surrogate networks for estimating empirical formulae discussed in section 2.2. Furthermore, we transfer the network parameters from different source models (trained on segment $0-60$ sec) which also adds to the overall uncertainty. In the following sections, we present results for clean and noisy data conditions. A summary of the mean and standard deviation estimates for the unknown parameters under different test conditions is presented in table 7.
@@
Table 7: Predicted unknown parameters (multi-stage Transfer Learning). The mean and standard deviation of the predicted unknown parameters when predicted using multi-stage transfer learning with clean data and Gaussian noisy data. The mask function is applied to the trainable variable to ensure positive values for predictions. Furthermore, scaling factors are used to convert engine parameter estimates from neural networks back to their original scales. For clean data, we show results obtained by using 151 and 301 data points while for noisy data, we show results obtained by using 301 data points. With noisy data, more data points are required for accurate parameter and state estimations.
Duration (sec) $A_{egrmax}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$ Remarks
Parameter estimates, Clean data
$0\sim 60$ $1.009\pm 0.0017$ $1.022\pm 0.0003$ $1.004\pm 0.0008$ $1.003\pm 0.0035$ 151 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.922\pm 0.0487$ $1.0185\pm 0.0036$ $0.977\pm 0.0183$ $0.850\pm 0.0967$
$120\sim 180$ $0.955\pm 0.0117$ $0.991\pm 0.0036$ $1.012\pm 0.0017$ $1.013\pm 0.0041$
$180\sim 240$ $1.007\pm 0.0017$ $1.019\pm 0.0004$ $1.003\pm 0.0025$ $0.992\pm 0.0072$
$240\sim 300$ $0.944\pm 0.0980$ $0.999\pm 0.0073$ $1.001\pm 0.0025$ $0.988\pm 0.0103$
$0\sim 60$ $1.004\pm 0.0010$ $1.024\pm 0.0003$ $1.007\pm 0.000$ $1.005\pm 0.0012$ 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.971\pm 0.0470$ $1.012\pm 0.0046$ $1.000\pm 0.0833$ $0.980\pm 0.0561$
$120\sim 180$ $0.964\pm 0.0123$ $0.992\pm 0.0011$ $1.008\pm 0.0025$ $1.007\pm 0.0056$
$180\sim 240$ $1.004\pm 0.0013$ $1.019\pm 0.0003$ $1.003\pm 0.0017$ $0.992\pm 0.0054$
$240\sim 300$ $0.900\pm 0.1033$ $1.006\pm 0.0022$ $0.998\pm 0.0042$ $0.984\pm 0.0137$
Parameter estimates, Noisy data.
@@
$0\sim 60$ $0.997\pm 0.1103$ $1.021\pm 0.0061$ $0.999\pm 0.0125$ $0.999\pm 0.0452$ 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.952\pm 0.0693$ $1.011\pm 0.0072$ $1.007\pm 0.0108$ $1.021\pm 0.0675$
$120\sim 180$ $0.803\pm 0.1047$ $0.974\pm 0.0148$ $0.998\pm 0.0125$ $0.984\pm 0.0734$
$180\sim 240$ $1.004\pm 0.0353$ $1.019\pm 0.0051$ $0.993\pm 0.0117$ $0.978\pm 0.0271$
$240\sim 300$ $0.793\pm 0.1333$ $1.002\pm 0.0103$ $0.997\pm 0.0117$ $0.985\pm 0.0328$
True value $1.000$ $1.000$ $1.000$ $1.000$ Normalized values for confidentiality
4.1 Parameter and state estimation under clean conditions
The estimates for the unknown parameters $\mathbf{\Lambda}$ obtained by training the hybrid PINN model for each data segment with the multi-stage transfer learning approach are illustrated in figure 6 and B.1.1. The variability in parameter estimation is obtained by computing the expected value and variance across thirty independent runs for each segment. The estimate of unknown parameters $\mathbf{\Lambda}$ represent the values obtained at the end of 35,000 epochs. We present a comparison between models trained with 151 and 301 residual points to validate our assertion that 151 points are adequate for accurate parameter estimation in clean data. Training with fewer data improves computational efficiency and hence, is preferred for field applications. From the violin plots in figure 6 and B.1.1, we note a high variation in all parameter estimates at t = 1-2 min. This is expected, as in this segment, both dynamic and idling conditions are encountered which pose challenges for engine state estimation using neural networks due to spectral bias [38]. With 301 data points, we observe a sightly reduced variance of parameter estimation across all $\mathbf{\Lambda}$ but with increased computation cost. This trade-off between requirements of computational time and accuracy needs to be considered for field application.
Figure 7 illustrates the prediction results for the eight output states defined in equations 3 through 10. Due to confidentiality, the dynamics shown in this figure are normalized on a scale 0-1 using the min-max values as:
@@ 4.1 Parameter and state estimation under clean conditions
where $x_{\text{max}}$, $x_{\text{min}}$ maximum and minimum value of the true data of $x$ variable within the time duration considered. Here, it is noteworthy that states $\tilde{u}_{egr1},\tilde{u}_{egr2}$, and $\tilde{u}_{vgt}$ are predicted directly using the pre-trained DeepONet model. The PINN model prediction indicates good match with simulated ground truth data for states $p_{im},p_{em}$, and $\omega_{t}$. However, states $T_{1}$ and $x_{r}$, which are approximated by the PINN model, exhibit higher standard deviations from simulated ground truth data. This can be attributed to the relatively smaller influence of these two state variables on gas dynamics. Prediction for these two states depend solely on the gradient of the physics-based loss term which is small leading to higher variation in state estimation due to numerical error. Figure 8 shows the different gas flow rates obtained as intermediate states by using relationships between primary dynamics and other quantities of interest (QoI). Overall, the prediction from PINN model demonstrates good approximation to ground truth values for all states. We note that state $A_{egr}$ determines the area of EGR valve that is controlled by the EGR actuator. We observe that the prediction from PINN model for segments 1-2 min and 4-5 min show higher deviation as compared to true values. Since this state estimate is directly related to the estimate of the unknown parameter $A_{egrmax}$, we expect the parameter estimate to exhibit high variability in these segments as well. We observe this behavior in figure 6 and B.1.1, which depict higher variation for the parameter $A_{egrmax}$ in these segments. We show additional ensemble plots by combining all 30 runs across the five segments in appendix B section B.1. The error band in these ensemble plots help estimate the epistemic uncertainty from this transfer learning model. We note higher variation in prediction of states $T_{1}$ and $x_{r}$, that correlates with our discussion about these states earlier in this section. Figure 19(b) shows a convergence plot for the four unknown parameters of interest for this study along with the range of expected values for time segment 3-4 minute. The uncertainty band shown in this plot is obtained by calculating point wise minimum and maximum values of these parameters across thirty independent runs. Note that all the four parameters converge around the true value of these parameters across all runs, which demonstrates the repeatability of this hybrid transfer learning framework.
@@
Fig. 6: Violin plot, multi-stage transfer learning, clean data, 151 data points showing the variation in unknown parameter predictions over 30 runs with clean data and 151 data points. 151 data points are considered for each of the known parameters (for each minute) under clean conditions. The violin plots are limited to the maximum and minimum values of the predicted parameters. The dots on the violin plots show a swarm plot indicating each predicted value for the 30 independent runs. The grey band shown indicates ± 7.5% error band. The dotted line in the centre indicates the true value of the unknown parameter. The top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
Fig. 7: Predicted states, multi-stage transfer learning, clean data, 151 data points. Primary gas flow dynamics states from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where clean data is used with 151 data points for online training. Note that states $\tilde{u}_{egr1},\tilde{u}_{egr2},$ and $\tilde{u}_{vgt}$ are outputs obtained from pre-trained DeepONet and are not trained during the transfer learning process.
Fig. 8: Gas flow rates, multi-stage transfer learning, clean data, 151 data points. Gas flow rates from transfer learning model (data normalized due to confidentiality). The plot shows a representative case where clean data is used with 151 data points for online training. The gas flow rates are used for parameter estimation through flow equations defined in appendix A.
4.2 Parameter and state estimation under noisy conditions
@@
To simulate a practical scenario, we assess the performance of our model when trained using noisy data. During our testing, we identified the need to use more data points to enhance the prediction accuracy of the transfer learning model. Consequently, we utilize 301 data points for training the PINN models under noisy conditions. Figure 9 illustrates the distribution of unknown parameters $\mathbf{\Lambda}$ over the five segments with thirty independent runs for each segment. With noise, we see higher variation across all parameter estimates which is expected due to the aleatoric uncertainty. Notably, amongst the four parameters, $A_{egrmax}$ demonstrates the largest increase in variation across all segments. In figure 10 and 11, we plot the prediction of primary gas dynamics and gas flow rates respectively for one sample run. We observe higher error in the prediction for state $A_{egr}$ when compared to its ground truth, correlating with the higher variability of $A_{egrmax}$ observed. Additional ensemble plots showing the predicted values and expected range of predictions for different QoIs are shown in Appendix B section B.2.
Fig. 9: Violin plot, multi-stage transfer learning, noisy data, 301 data points: Variation in engine parameter prediction over 30 runs with Gaussian noise in data signals with 301 data points. The dots on the violin plots shows a swarm plot indicating each predicted value for the 30 independent runs. The grey band shown indicates ± 7.5% error band. The dotted line in center indicates the true value of the unknown parameter. Higher variation in parameter estimation is observed across all time segments and represents the aleatoric uncertainty expected in field due to sensory noise.
Fig. 10: Predicted states, multi-stage transfer learning, noisy data, 301 data points: Prediction results for primary gas flow dynamics states from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where 301 noisy data points are used for online training of hybrid PINN model.
@@
Fig. 11: Gas flow rates, multi-stage transfer learning, noisy data, 301 data points: Gas flow rates from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where 301 noisy data points are used for online training.
@@ 5.1 Training of multi-head neural networks
As discussed earlier, few-shot transfer learning involves two Phases: (1) offline training of the neural network with multiple heads and (2) training the head and unknown parameters for the inverse problem. The training in Phase I is done with 240 heads for each variable; each head is used for estimating a one-minute data segment. The input to the network is $0-60$ sec ($dt=0.2$ sec) normalized to $[-1,1]$. The first 120 heads correspond to the state variable for Case-I, and the second 120 heads correspond to the state variable for Case-IV (ref Table 5). We have shown the network detail and training error for each state variable in Table 8. We trained these networks using the Adam optimizer with an MSE loss function between the labelled and the predicted output. In the case of $N_{1}(t,\bm{\theta})$, as there are two outputs, an appropriate weight is considered while summing the two losses. The weighs of the network are initialized from a Gaussian distribution with mean zero and standard deviation $\sigma=\sqrt{2/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$ except for the last layer. The last layer is initialized with a standard deviation of $\sigma=\sqrt{10/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$. This shows a better accuracy compared to an initialization with a similar standard deviation as the hidden layers. This may be because, as $\text{dim}_{\text{out}}$ in the last layer has more neurons, the standard deviation becomes smaller compared to the previous layers. Furthermore, to check the applicability of the method for practical application, the training in Phase-I is done with different parameters and the transfer learning (Phase-II) is done with different parameters, as discussed in section 2.1 (ref Table 1).
Table 8: Training of multi-head networks for few-shot transfer learning (Phase-I): Multi-head neural networks considered to approximate the state variables and $x_{r}$ and $T_{1}$. The input to the network is time ($t$). The input time is 0 to 60 sec with $\delta t=0.2$ sec, and the input is scaled between $[-1,1]$. Each head approximates a minute of the variable. We considered 240 heads for each output approximating 240 minutes of dynamics. Thus, network $\mathcal{N}_{2}(t,\bm{\theta}_{2})$, $\mathcal{N}_{3}(t,\bm{\theta}_{3})$ and $\mathcal{N}_{4}(t,\bm{\theta}_{4})$ has 240 output, however network $\mathcal{N}_{1}(t,\bm{\theta}_{1})$ has 480 output as it approximate two ($p_{em}$ and $p_{im}$) variables. The ”Output transformation” and scaling are considered as per Table 1. The $L_{2}$ error shows % relative $L_{2}$ error in training for each variable.
Network $\mathcal{N}_{1}(t,\bm{\theta}_{1})$ $\mathcal{N}_{2}(t,\bm{\theta}_{2})$ $\mathcal{N}_{3}(t,\bm{\theta}_{3})$ $\mathcal{N}_{4}(t,\bm{\theta}_{4})$
Output $p_{em}$, $p_{im}$ $x_{r}$ $T_{1}$ $\omega_{t}$

### v2

@@ 3 Results: Multi-stage Transfer Learning
In this section, we present results for the inverse problem solved using the hybrid multi-stage TL approach outlined in section 2.5. We consider two scenarios: i) training with clean data, and ii) training with noisy data. Since we use simulated data for model training, we added Gaussian noise to these simulated outputs to mimic real-world conditions. Based on expected noise from field sensors, we added Gaussian noise values of 3%, 3%, 1%, 5% and 3% to simulated outputs for $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ respectively. The input signal required for the Simulink model viz. control input $\bm{u}=\{u_{\delta},u_{egr},u_{vgt}\}$ and engine speed $n_{e}$, are obtained from field measurements on an engine test bed. As discussed in section 2.5, we train the hybrid PINN model on individual one minute data segments. In the first segment, the complete hybrid PINN model is trained. Complete training of this hybrid PINN model is computationally expensive since it requires learning all network parameters and requires more training epochs. To reduce computational time for subsequent time segments, we transfer the network weights and biases to the multi-stage TL model.
@@
We approximate $p_{im}$, $p_{em}$, $\omega_{t}$, $T_{1}$, and $x_{r}$ using neural networks (as shown in Table 4) and incorporate mask function and scaling for the unknown parameters, as shown in Table 1. We note that output transformation, mask function, and scaling are important to ensure the network outputs and unknown parameters lie in $\mathbb{R^{+}}$. Additionally, we incorporate self-adaptive weights for the data loss term of $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$, initialized from a Gaussian random variable with mean $\dfrac{\alpha X_{data}}{\max(X_{data})}$, $X=\{p_{em},p_{im},\omega_{t},W_{egr}\}$ and standard deviation $\sigma=0.1$, where $\alpha$ is a heuristically determined weight multiplier ($10^{3}$, $10^{3}$, $10^{3}$, and $2\times 10^{3}$ for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ respectively). The self-adaptive weight $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ are initialized from a Gaussian random variable with mean $1500$ and standard deviation 0.1. A softplus masking function is applied to the self-adaptive weights to ensure that they remain positive. The unknown parameters for the first 0-60 sec are initialized from a Gaussian random variable with standard deviation $\sigma=0.2$ and mean near the true values of the parameters. For calculating PINN loss defined in Eq. (18), 301 residual points are used for individual PINN loss terms. The parameters of the neural network, unknown engine parameters, and self-adaptive weights are optimized using the Adam optimizer [41] with piece-wise constant learning rates for each stage with multiple optimizers. Stage I extends from 0-20K epochs, Stage II between 20-30K epochs, and Stage III from 30-35K epochs (ref Fig. 3 for further details). To estimate the epistemic uncertainty, we train the hybrid PINN model separately for each data segment to generate an ensemble of 30 runs. This uncertainty arises from the random initialization of self-adaptive weights $\mathbf{\lambda}$, unknown parameter values $\mathbf{\Lambda}$, and dropouts included in the surrogate networks for estimating empirical formulae discussed in section 2.2. Furthermore, we transfer the network parameters from different source models (trained on segment $0\sim 60$ sec) which also adds to the overall uncertainty. In the following sections, we present results for clean and noisy data conditions. A summary of the mean and standard deviation estimates for the unknown parameters under different test conditions is presented in Table 6.
@@
Table 6: Predicted unknown parameters (multi-stage Transfer Learning). The mean and standard deviation of the predicted unknown parameters with multi-stage transfer learning using clean data and Gaussian noisy data. For clean data, we show results obtained by using 151 and 301 data points while for noisy data, we show results obtained by using 301 data points. With noisy data, more data points are required for accurate parameter and state estimations.
DURATION (min) $A_{egrmax}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$
Clean data, 151 data points for each known variables (for each 60 sec.).
$0\sim 1$ $1.009\pm 0.0017$ $1.022\pm 0.0003$ $1.004\pm 0.0008$ $1.003\pm 0.0035$
$1\sim 2$ $0.922\pm 0.0487$ $1.0185\pm 0.0036$ $0.977\pm 0.0183$ $0.850\pm 0.0967$
$2\sim 3$ $0.955\pm 0.0117$ $0.991\pm 0.0036$ $1.012\pm 0.0017$ $1.013\pm 0.0041$
$3\sim 4$ $1.007\pm 0.0017$ $1.019\pm 0.0004$ $1.003\pm 0.0025$ $0.992\pm 0.0072$
$4\sim 5$ $0.944\pm 0.0980$ $0.999\pm 0.0073$ $1.001\pm 0.0025$ $0.988\pm 0.0103$
Clean data, 301 data points for each known variables (for each 60 sec.).
$0\sim 1$ $1.004\pm 0.0010$ $1.024\pm 0.0003$ $1.007\pm 0.000$ $1.005\pm 0.0012$
$1\sim 2$ $0.971\pm 0.0470$ $1.012\pm 0.0046$ $1.000\pm 0.0833$ $0.980\pm 0.0561$
$2\sim 3$ $0.964\pm 0.0123$ $0.992\pm 0.0011$ $1.008\pm 0.0025$ $1.007\pm 0.0056$
$3\sim 4$ $1.004\pm 0.0013$ $1.019\pm 0.0003$ $1.003\pm 0.0017$ $0.992\pm 0.0054$
$4\sim 5$ $0.900\pm 0.1033$ $1.006\pm 0.0022$ $0.998\pm 0.0042$ $0.984\pm 0.0137$
Noisy data, 301 data points for each known variables (for each 60 sec.).
@@
$0\sim 1$ $0.997\pm 0.1103$ $1.021\pm 0.0061$ $0.999\pm 0.0125$ $0.999\pm 0.0452$
$1\sim 2$ $0.952\pm 0.0693$ $1.011\pm 0.0072$ $1.007\pm 0.0108$ $1.021\pm 0.0675$
$2\sim 3$ $0.803\pm 0.1047$ $0.974\pm 0.0148$ $0.998\pm 0.0125$ $0.984\pm 0.0734$
$3\sim 4$ $1.004\pm 0.0353$ $1.019\pm 0.0051$ $0.993\pm 0.0117$ $0.978\pm 0.0271$
$4\sim 5$ $0.793\pm 0.1333$ $1.002\pm 0.0103$ $0.997\pm 0.0117$ $0.985\pm 0.0328$
True value (normalized)
$1.000$ $1.000$ $1.000$ $1.000$
3.1 Parameter and gas flow dynamics estimation
We evaluate the performance of our multi-stage TL model by comparing the state and parameter estimates against known ground truth values by simulating online testing conditions. Due to confidentiality, the dynamics shown in the results are normalized on a scale 0-1 using their min-max values as:
@@ 3.1 Parameter and gas flow dynamics estimation
The variability in parameter estimation is obtained by computing the expected value and variance across thirty independent runs for each segment. The estimate of unknown parameters $\mathbf{\Lambda}$ represent the values obtained at the end of 35,000 epochs. For clean data, we compared the effect of number of data points on parameter estimation by evaluating our model with 151 and 301 points. We observe that for clean data, 151 data points are sufficient for accurate estimation of parameters and gas flow states. See Fig.5 showing the distribution of estimated parameters for 30 independent runs. Higher variation is observed in time segment 1-2 minute for all parameters likely due the engine entering idling condition from a dynamical state which induces bias in PINN training With 301 data points, we observe a sightly reduced variance of parameter estimation across all $\mathbf{\Lambda}$ (see Fig. A-3). Training with fewer data improves computational efficiency and hence, is preferred for field applications. This trade-off between requirements of computational time and accuracy needs to be considered for field application. Overall, the mean estimated values are reasonably close to the ground truth values.
@@
Fig. 5: Violin plot, multi-stage TL, clean data, 151 data points: variation in unknown parameter predictions over 30 runs with clean data and 151 data points. 151 data points are considered for each of the known parameters (for each minute) under clean conditions. The violin plots are limited to the maximum and minimum values of the predicted parameters. The dots on the violin plots show a swarm plot indicating each predicted value for the 30 independent runs. The grey band shown indicates ± 7.5% error band. The dotted line in the centre indicates the true value of the unknown parameter. The top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
@@
To simulate a practical scenario, we assess the performance of our model when trained using noisy data. During our testing, we identified the need to use more data points to enhance the prediction accuracy of the TL model. Consequently, we utilize 301 data points for training the PINN models under noisy conditions. From Fig. 6, we observe that predictions from PINN model show a good match with ground truth data for states $p_{im},p_{em}$, and $\omega_{t}$ under noisy conditions. However, states $T_{1}$ and $x_{r}$ exhibit higher standard deviations from simulated ground truth data. This can be attributed to the relatively smaller influence of these two state variables on gas dynamics. Prediction for these two states depend solely on the gradient of the physics-based loss term which is small leading to higher variation in state estimation due to numerical error. Overall, the prediction from PINN model demonstrates good approximation to ground truth values for all states. We observe from the gas flow rate predictions shown in Fig. 7 that the prediction for state $A_{egr}$ from PINN model shows higher deviation from expected ground truth. Since this state estimate is directly related to the estimate of the unknown parameter $A_{egrmax}$, we expect the parameter estimate to exhibit high variability. This can be observed in Fig. 8 which illustrates the distribution of unknown parameters $\mathbf{\Lambda}$ over the five time segments with thirty independent runs for each segment. With noise, we see higher variation across all parameter estimates which is expected due to the aleatoric uncertainty. Additional ensemble plots showing the predicted values and expected range of predictions for different QoIs are shown in Appendix section F.
Fig. 6: Engine states, multi-stage TL, noisy data, 301 data points: Comparison between hybrid-PINN prediction and ground truth. The model shows reasonable accuracy when predicting engine states considering the rapidly changing engine dynamics across the five minute long test signal.
@@
Fig. 7: Gas flow rates, multi-stage TL, noisy data, 301 data points: Gas flow rates from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where 301 noisy data points are used for online training.
Fig. 8: Violin plot, multi-stage TL, noisy data, 301 data points: 301 data points are considered for each of the known parameters (for each minute) under noisy conditions. The violin plots are limited to the maximum and minimum values of the predicted parameters. The results are shown after normalizing the true value of the unknown parameters.
@@ 4.1 Training of multi-head neural networks
As discussed earlier, few-shot TL involves two Phases: (1) offline training of the neural network with multiple heads and (2) training the head and unknown parameters for the inverse problem. The training in Phase I is done with 240 heads for each variable; each head is used for estimating a one-minute data segment. The input to the network is $0\sim 60$ sec ($dt=0.2$ sec) normalized to [1,1][-1,1]. The first 120 heads correspond to the state variable for Case-I dataset, and the second 120 heads correspond to the state variable for Case-IV dataset (ref Table 5). We have considered four NN similar to those in Table 4 with the same transformation and scaling. However, with different sizes and with multiple heads. The training error for each state variable and $x_{r}$ and $T_{1}$ are in Table A-3 (in Appendix G). We trained these networks using the Adam optimizer with an MSE loss function between the labelled and the predicted output. In the case of $N_{1}(t,\bm{\theta})$, as there are two output state variables, an appropriate weight is considered while summing the two losses. The weighs of the network are initialized from a Gaussian distribution with mean zero and standard deviation $\sigma=\sqrt{2/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$ except for the last layer. The last layer is initialized with a standard deviation of $\sigma=\sqrt{10/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$. This shows a better accuracy compared to an initialization with a similar standard deviation as the hidden layers. This may be because, as $\text{dim}_{\text{out}}$ in the last layer has more neurons, the standard deviation becomes smaller compared to the previous layers. Furthermore, to check the applicability of the method for practical application, the training in Phase-I is done with different parameters and the TL (Phase-II) is done with different parameters, as discussed in section 2.1 (ref Table 1). We note that in the case of few-shot transfer learning, the PINN network sizes are comparatively larger than the multi-stage network sizes as they approximate multiple heads. However, since only the last layer is trained and gradients are calculated outside the training loop, this framework is computationally more efficient.
4.2 Parameter and gas flow dynamics estimation
After training,, the next step is to perform PINN to identify the engine parameters (Phase-II). The output layer ($\bm{W}_{out}$, $b_{out}$) of the multi-head networks is replaced with one output and trained for a new control vector ($\bm{u}=\{u_{\delta},u_{vgt},u_{egr}\}$) and engine speed ($n_{e}$). The weights of the output layer are initialized from a random Gaussian distribution with mean zero and standard deviation $\sigma=\sqrt{2/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$. The weights are sorted after initialization, and it shows better and faster convergence. The biases are initialized with zero. Similar to the PINN and TL method, each of the self-adaptive weights for the data loss term of $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ are initialized from a Gaussian random variable with mean $\dfrac{\alpha X_{data}}{\max(X_{data})}$, $X=\{p_{em},p_{im},\omega_{t},W_{egr}\}$ and standard deviation $\sigma=0.1$, where $\alpha$ is a initial weight given ($10^{3}$, $10^{3}$, $10^{3}$ and $2\times 10^{3}$ for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ respectively). The self-adaptive weight $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ are initialized from a Gaussian random variable with mean $20$ and standard deviation 0.1. We consider a softplus mask function to the self-adaptive weights so that the weights are always positive. The unknown parameters are initialized from a Gaussian random variable with standard deviation $\sigma=0.2$ and mean near the true values of the parameter. Furthermore, as these quantities are always positive, we consider mask functions and scaling as shown in Table 1. Also note that the trainable parameters of the problem are the weights and bias of the last layer of the neural networks, the unknown parameters of the engine and the self-adaptive weights. These are optimized using eight different Adam optimizers: one for the weights and the biases of the neural networks, one for the unknown parameters, and one for each of the six self-adaptive weights. The loss function is given by the Eq (18) as discussed in section 2. We would also like to mention that we only optimized the self-adaptive weights up to $20\times 10^{3}$ epoch. After that, we fixed the self-adaptive weights with the value at $20\times 10^{3}$ epochs. We consider total $30\times 10^{3}$ epochs. Furthermore, as discussed earlier, in order to reduce the computational cost, we perform the derivative of the output with respect to the input outside the training loop as follows,
$y=S_{p}\left(\bm{h}(t)\bm{W}_{out}+b_{out}\right)$ (22a)
$\begin{split}\dfrac{dy}{dt}=&\dfrac{1}{1+\exp[\bm{h}(t)\bm{W}_{out}+b_{out}]}\times\\ &\exp\left[\bm{h}(t)\bm{W}_{out}+b_{out}\right]\dfrac{d\bm{h}(t)}{dt}\bm{W}_{out}\end{split}$ (22b)
where $S_{p}$ is softplus function (refer Table A-3), $\bm{h}(t)$ is the output of the last hidden layer, $\bm{W}_{out}$ and $b_{out}$ are the weights and bias of the output layer, $\dfrac{d\bm{h}(t)}{dt}$ is the derivative of the output of the last hidden layer with respect to input $t$. From Eq. (22b), the derivative term is equal to $\dfrac{d\bm{h}(t)}{dt}$, and as the hidden layers are not trained, this remains fixed and can be evaluated outside the network training loop. The size of $\dfrac{d\bm{h}(t)}{dt}$ is a matrix with size $n_{t}\times n_{h}$, where $n_{t}$ is the number of input points in time, and $n_{h}$ is the number of neurons in the last hidden layer.
We consider two different scenarios with clean data without additional noise in the field data (simulated). However, field data may be noisy in practice due to the inherent instrument noise or other reasons. Thus, to simulate a more practical scenario and the applicability of the proposed method, we consider added noise and perform PINN. We added a Gaussian noise of 3%, 3%, 1%, 5% and 3% added to the simulated data of $p_{em}$, $p_{im}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ respectively. We study five minutes of engine operation and sub-divide into five independent problems, each studied independently. For each one minute, we consider 301 residual points with equal $dt=0.2$ sec. For the first case of clean data, we consider $151$ data points ($dt=0.4$ sec.) for the data loss of $p_{em}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$, and for the second case of clean data we consider $301$ data points. In the case of noisy data, we consider 301 data points. The size of the self adaptive weights for $\lambda_{p_{em}}$, $\lambda_{p_{im}}$, $\lambda_{\omega_{t}}$ and $\lambda_{W_{egr}}$ is equal to the number of data point considered, and the size self-adaptive weight for $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ is $1\times 1$. Before discussing further, we would like to mention that, there are three different types of uncertainties: (i) uncertainty associated with the initialization of the last layer of the neural network and other parameters (e.g. unknown parameters, self-adaptive weights), (ii) uncertainty in the pre-trained network due to the dropout considered, (iii) uncertainty due to the added noise in the field variables (in case of noisy data). In order to check the effect of initialization and randomness on the accuracy, we consider $30$ independent runs for each one-minute data segment.